OpenStudy (tanner23456):

y = 1 + 6x^(3/2) ; 0<= x <= 1 Find the Length

OpenStudy (tanner23456):

I know I'm supposed to use \[L=∫\sqrt{1+(dy/dx)^2}dx\]

OpenStudy (anonymous):

right so can you get the derivative first

OpenStudy (anonymous):

dy/dx = 0 +9*x^(1/2) (dy/dx)^2 = 81*x this is what you should be able to get first

OpenStudy (tanner23456):

Right, that's what I got so it's: \[\int\limits_{0}^{1} \sqrt{1+81x} dx\] or \[\int\limits_{0}^{1}(1+81x)^{1/2} dx\] Where to I go from here?

OpenStudy (anonymous):

so far no problem, the last part is to intergrate

OpenStudy (anonymous):

\[2/243(1+81x)^{3/2}|^1_0\]

OpenStudy (tanner23456):

Yeah, for some reason I just completely forgot about u substitution. Thanks for your help.